Consider the surface S=(x,y,z)∈R3|x2+y2+z2=9.

(a)Define an R3−R function f such that S is the contour surface of f at level 9.

(b)Find an equation for the plane V that is tangent to S at the point (x,y,z) = (2,1,2).

(c) Sketch the surface Sin R3,together with a section of the plane V to illustrate that V is tangent to S at the point (2,1,2).


1
Expert's answer
2021-06-01T14:35:32-0400

Solution :-

"s=\{(x,y,z)\in R^3 |x^2+y^2+z^2=9 \}"

(a) "f:R^3 \rightarrow R \\ such \ that \ S \ is \ the \ contour \ surface \ of \ f \ at \ level \ 9 \\
so \ \ \ \ \ f(x,y,z)=x^2+y^2+z^2-9=0"


(b)"f(x,y,z)=x^2+y^2+z^2-9 \\"

than "df=2xdx+2ydy+2zdz"

"df|_{(2,1,2)}=4\hat{i}+2\hat{j}+4\hat{k}"

so equation of tangent plane V will be

4(x-2)+2(y-1)+4(z-2)=0

"\implies 4x-8+2y-2+4z-8=0"

"\implies \boxed{4x+2y+4z=18}"


(c)